# Converging to Stability in Two-Sided Bandits: The Case of Unknown Preferences on Both Sides of a Matching Market

### v2

@@ 1. Introduction
The classic literature on two-sided matching (Gale and Shapley, 1962; Roth and Xing, 1997; Haeringer and Wooders, 2011, e.g.), encompassing applications including long- and short-term labor markets, dating and marriage, school choice, and more, has typically focused on situations where agents are aware of their own preferences. The problem of learning preferences while participating in a repeated matching market first started receiving attention in the AI literature in the work of Das and Kamenica (2005), and the general idea of two-sided matching under unknown preferences has since been studied in economics and operations research as well  (Lee and Schwarz, 2009; Johari et al., 2022). This area of research has received renewed attention in the last few years, along with novel theoretical insights into convergence properties of upper-confidence-bound (UCB) style algorithms (Liu et al., 2021; Kong et al., 2022; Zhang et al., 2022).
@@
The two-sided matching problem involves agents on two sides of a market who have preferences for each other but cannot communicate explicitly. The goal is to create a matching process that ensures stability, where no pairs of agents would rather be matched with each other over their current match. The existence of such matchings was famously demonstrated constructively in the Gale-Shapley algorithm (Gale and Shapley, 1962) which structured around one side of the market proposing and the other side choosing whether to accept proposals. This theory has been applied to various markets, such as matching medical students to residencies (Roth and Peranson, 1999) and students to schools (Abdulkadiroğlu et al., 2005), with the assumption that agents know their own preferences. Considerable interest has also been shown in the AI community regarding two-sided matching in the presence of various constraints, such as diversity constraints (Aziz et al., 2022; Benabbou et al., 2018). A version of the problem, called two-sided bandits, was introduced in Das and Kamenica (2005), where agents engage in repeated matching without prior knowledge of preferences, and the impact of the matching mechanism on convergence to stability was studied. Recent work has also focused on computing the probability that a matching is stable or finding the matching with the highest probability of being stable under different models of preference uncertainty (Aziz et al., 2020).
@@
The exploration-exploitation dilemma that characterizes bandit problems is further complicated in two-sided bandits by uncertainty on not just the values of the “arms” of the bandit, but also uncertainty on whether the arms will accept or reject your attempt to pull them (your proposal). This additional uncertainty arises from the multi-agent nature of the problem; instead of there being just one player, there are multiple players competing for the arms, and the arms also have preferences associated with the players. A popular solution choice for this type of problem involves the explore-then-commit style of algorithms (Pagare and Ghosh, 2023; Zhang and Fang, 2024). This flavor of solution includes a first phase of “learning” preferences, and a second phase using what has been learned in the prior one to commit to a choice. This involves spending some dedicated time trying to figure out the preferences of the agents. In this paper we are more interested in approaches where learning happens simultaneously with the matching process. One such algorithm, CA-UCB (Liu et al., 2021), assumes that at each round all the agents on the receiving side (henceforth arms) get to view all their proposals before deciding which one to accept. The algorithm provably converges to stability under the assumptions that (1) all arms have complete knowledge of their preferences, and (2) these preferences are common knowledge, so the proposers (henceforth players) also know them. The algorithm is a variation of the well-known family of UCB algorithms.
Contributions
@@
In this paper, we are interested in the more general case, where neither the players nor the arms start off with knowledge of their own preferences. This problem is significantly more complex because it adds a layer of uncertainty that could potentially lead to agents converging to an unstable equilibrium more easily than in the case with one-sided uncertainty and common knowledge of the other side’s preferences. We tackle this in two steps. First, we relax the common knowledge assumption and assume only that arms are aware of their own preferences.
Recent works have used frameworks where players communicate with each other to find a matching in this setting (Zhang et al., 2022; Zhang and Fang, 2024). While there are approaches which allow for and design explicit communication frameworks (Zhang et al., 2022), and also others where players only communicate indirectly to the extent of carefully orchestrating “collisions” (Zhang and Fang, 2024), we resolve the question of convergence in the more traditional non-cooperative setting of stable matching when arm preferences are not common knowledge. Our algorithm uses the concept of “plausible sets” (arms that an player might be able to “win” in the next time step) from the CA-UCB algorithm (Liu et al., 2021). We show how to define and update these sets in order to develop a provably convergent algorithm (which we call Optimistic CA-UCB or OCA-UCB). The case with two-sided uncertainty is more challenging. We introduce the Probabilistic Conflict Avoiding - Simultaneous Choice Algorithm (PCA-SCA), a provably convergent framework that leverages an additional optimistic estimate for the probability of winning each contentious situation. Although we prove convergence using UCB-style reward estimates, the framework accommodates other methods such as Thompson Sampling. Experimentally, we demonstrate convergence to stability under both UCB and Thompson-based reward estimates and evaluate how the convergence speed scales with the number of agents and the degree of preference heterogeneity.
2. Setting
@@
There are $N$ proposers (henceforth players) and $K$ proposees (henceforth arms). To ensure that each player can be matched, we assume that $N\leq K$ (Liu et al., 2020, 2021; Basu et al., 2021; Sankararaman et al., 2021). The set of players is denoted by $\left\{p_{i}\right\}_{i=1}^{N}$ and the set of arms by $\left\{a_{k}\right\}_{k=1}^{K}$. For each $p_{i}$, there is a distinct (unknown) mean reward $\mu_{i}^{k}$ associated with each arm $a_{k}$ with unit variance. If $\mu_{i}^{k}>\mu_{i}^{j}$, we say that $p_{i}$ prefers $a_{k}$ over $a_{j}$ denoted by $a_{k}\succ_{p_{i}}a_{j}$. Similarly, the arms also have distinct mean rewards associated with each player. The arms may or may not have this information available to them a priori depending on the setting in consideration.
@@
We use the classic notion of a stable matching: a bipartite matching from the group of players to the group of arms such that no player and arm prefer each other over their current matching i.e $\nexists(p_{i},m_{i}(t))$ such that $p_{i}$ prefers an arm $a_{j}\neq m_{i}(t)$ and $a_{j}$ also prefers $p_{i}$ over its current match. There may be more than one stable match. Our primary focus in this work is on convergence to stability. Another quantity we study is maximum Player Pessimal Regret. We use the Gale-Shapley algorithm with arms as proposers in order to calculate the player-pessimal stable match (Gale and Shapley, 1962). Denote the reward received by player $p_{i}$ in a player pessimal match by $X_{i}$. Then, we define maximum player-regret at each time-step as $R(t)=\underset{i\in N}{\max}\left[X_{i}-X_{i}^{m(t)}\right]$ which is the maximum difference (over players) between the reward a player received in its current match and the reward it would have received in a stable matching that player-pessimal.
@@
The APCK model has been studied and a decentralized solution approach is described in the Conflict Avoiding Upper Confidence Bound (CA-UCB) algorithm (Liu et al., 2021). Our main contribution is to detail decentralized approaches to solutions for Scenarios 1 and 2. We begin with a quick recap of CA-UCB in the APCK model.
3. APCK Model and CA-UCB
@@
In a recent paper, the APCK model (Scenario 0) was studied (Liu et al., 2020). Subsequently, a solution was proposed that is decentralized, the CA-UCB algorithm (Liu et al., 2021). CA-UCB finds a stable match in the APCK model by avoiding conflicts using the notion of a “plausible set.” The plausible set for a player $p_{i}$ at time-step $t$ is defined as $S_{i}(t):=\{a_{k}:p_{i}\succ_{a_{k}}p_{j},\text{ where }\bar{A}_{j}(t-1)=a_{k}\}$ i.e. the set of all arms $a_{k}$ such that at time-step $(t-1)$ another player $p_{j}\neq p_{i}$ successfully pulled the arm, but $p_{i}$ knows that $a_{k}$ prefers $p_{i}$ over $p_{j}$. This set also includes the arms the player successfully pulled (if any) and all the arms that were unmatched at the previous time-step. This makes it so that $p_{i}$ only attempts arms where there is a possibility of it successfully pulling the arm, avoiding conflicts. Then, $p_{i}$ attempts to pull the arm with the highest UCB value among $S_{i}(t)$ The UCB value tracked by $p_{i}$ for an arm $a_{k}$ is calculated as:
$\text{UCB}_{i}^{k}=\hat{\mu}_{i}^{k}+\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t)}}$
where $\hat{\mu}_{i}^{k}$ is the empirical mean reward tracked by $p_{i}$ for $a_{k}$ up to time $t$, and $n_{i}^{k}(t)$ is the total number of samples that is used to calculate that empirical mean.
4. APKP Model and OCA-UCB
Note that the CA-UCB approach works only when both players and arms have full knowledge of the arms’ preferences, which players use to construct $S_{i}(t)$. When players lack this information, they must learn their positions in the arms’ preference orderings and recreate plausible sets. Recent game-theoretic approaches address this scenario. For instance, (Etesami and Srikant, 2024) allows proposers to maintain and update probability distributions over potential arms based on feedback from previous rounds. Accepted proposals increase the likelihood of future proposals to that arm, while rejections decrease it. These methods guarantee convergence to a stable matching without centralized coordination or full knowledge of preferences, achieving logarithmic regret in hierarchical (Etesami and Srikant, 2024) markets. Similarly, some approaches use mood-based state variables—like “content,” “discontent,” and “watchful”—to govern each player’s strategy at each time-step (Shah et al., 2024). Based on utility feedback, proposers decide whether to stick with their current match or experiment with new proposals. This ensures that over repeated interactions, proposers learn their preferences and the system converges to the proposer-optimal stable match, even in fully decentralized and information-limited markets. We present a much simpler approach to the solution which requires “collisions” with other players at most once for the players to learn arm preferences. We call it OCA-UCB (Optimistic CA-UCB) which exploits the fact that the arms still have full knowledge of their own preferences in order to achieve this.
Following the general structure of the CA-UCB algorithm, all players initially set their UCB estimates for the arms to $\infty$. However, now in addition to the UCB estimates, each player will also keep track of their position in the arms’ preference orderings. This can be done simply by having each player maintain two sets: $O^{k_{h}}_{(i)}$ the set of all players that $p_{i}$ believes are ranked higher than itself in $a_{k}$’s preference ordering, and $O^{k_{l}}_{(i)}$ which is the set of all players that $p_{i}$ believes are ranked lower than itself. Initially, for each player $p_{i}$, these sets are initialized to $O^{k_{h}}_{(i)}=\emptyset$ and $O^{k_{l}}_{(i)}=\left\{p_{j}\right\}_{j=1}^{N}\setminus p_{i}$. Thus, each player starts with the optimistic belief that it is the most preferred by each of the arms.
Notice, when updating plausible sets, the only information that $p_{i}$ requires is its’ relative positioning in $a_{k}$’s preference with respect to $p_{j}$ if $\bar{A}_{j}(t-1)=a_{k}$. This can be accomplished using the two sets described above. Since each player starts out with an optimistic view of itself in the arms’ preference orderings, conflicts are almost inevitable. This is because according to each player, all the arms are in the plausible set, even when they ought not to be. Suppose conflict occurs between two players, say $p_{i}$ and $p_{j}$, for an arm $a_{k}$ at time-step $t$, and player $p_{j}$ wins the conflict, then $p_{i}$’s belief with respect to $p_{j}$ and $a_{k}$ gets updated. More precisely, $O^{k_{h}}_{(i)}=O^{k_{h}}_{(i)}\cup\{p_{j}\}$ and $O^{k_{l}}_{(i)}=O^{k_{l}}_{(i)}\setminus\{p_{j}\}$. At future time steps, $p_{i}$ can use this information when considering $a_{k}$ in its plausible set. Functionally the rest of the algorithm is the same as CA-UCB. The only difference is that, at the end of each time-step, the players make use of the matching information and update $O^{k_{h}}_{(i)}$ and $O^{k_{l}}_{(i)}$ accordingly. The full algorithm is shown in Algorithm 1.
Algorithm 1 Optimistic Conflict Avoiding - UCB (OCA-UCB)
1:Input: $\lambda$, $H$ 2:Output: A stable match 3:for $t\in\{1,\cdots,H\}$ do 4:     for $i\in\{1,\cdots,N\}$ do 5:         if $t==1$ then 6:              $A_{i}(t)\leftarrow a_{k},k\sim\text{Uniform}(\left[K\right])$ 7:         else 8:              $D_{i}(t)\sim Ber(\lambda)$ 9:              if $D_{i}(t)==0$ then 10:                  $A_{i}(t)\leftarrow$ GetBestArm($p_{i},t$) 11:              else 12:                  $A_{i}(t)\leftarrow A_{i}(t-1)$ 13:              end if 14:         end if 15:     end for 16:     for $k\in\{1,\cdots,K\}$ do 17:         $p_{w}\leftarrow$ ResolveRequests($a_{k},t$) 18:         if $p_{w}\neq\emptyset$ then 19:              $r_{p_{w}},r_{a_{k}}\leftarrow$ Sample-Rewards($p_{w},a_{k}$) 20:              Arm-Update-Belief($a_{k},p_{w},r_{a_{k}}$) 21:              Player-Update-Belief($p_{w},a_{k},r_{p_{w}}$) 22:              $L\leftarrow\left\{p_{j}:A_{j}(t)=a_{k}\right\}\setminus\left\{p_{w}\right\}$ 23:              for $p_{l}\in L$ do 24:                  UpdateConflictBelief($p_{l},p_{w},a_{k}$) 25:              end for 26:         end if 27:     end for 28:end for 1:procedure GetBestArm ($p_{i}$, $t$) 2:     $S_{i}(t)\leftarrow\emptyset$ 3:     for $k\in\{1,\cdots,K\}$ do 4:         $p_{\bar{p}}\leftarrow$ $p_{j}\text{ s.t. }\bar{A}_{j}(t-1)=a_{k}$ 5:         if $p_{\bar{p}}\in O_{(i)}^{k_{l}}\vee p_{\bar{p}}=\emptyset$ then 6:              $S_{i}(t)\leftarrow S_{i}(t)\cup\{a_{k}\}$ 7:         end if 8:     end for 9:     $k\leftarrow\underset{a_{j}\in S_{i}(t)}{\mathop{\mathrm{arg\,max}}\limits}\left[\text{UCB}_{i}^{j}\right]$ 10:     return $a_{k}$ 11:end procedure 1:procedure UpdateConflictBelief($p_{i},p_{j},a_{k}$) 2:     $O_{(i)}^{k_{l}}\leftarrow O_{(i)}^{k_{l}}\setminus p_{j}$ 3:     $O_{(i)}^{k_{h}}\leftarrow O_{(i)}^{k_{h}}\cup\{p_{j}\}$ 4$\triangleright$ $p_{i}$ lost conflict with $p_{j}$ for $a_{k}$ 5:end procedure
4.1. Proof of OCA-UCB Convergence
As Algorithm 1 progresses, each of the players will have complete sets that represent the arms’ true preferences with respect to the other players. In fact, we show theoretically that this approach to the solution does not affect the convergence guarantee of the CA-UCB algorithm in the APCK Model in Theorem 4.1.

### v4

@@ 1 Introduction
The classic literature on two-sided matching (Gale and Shapley, 1962; Roth and Xing, 1997; Haeringer and Wooders, 2011, e.g.), encompassing applications including long- and short-term labor markets, dating and marriage, school choice, and more, has typically focused on situations where agents are aware of their own preferences. The problem of learning preferences while participating in a repeated matching market first started receiving attention in the AI literature in the work of Das and Kamenica (2005), and the general idea of two-sided matching under unknown preferences has since been studied in economics and operations research as well  Lee and Schwarz (2009); Johari et al. (2022). This area of research has received renewed attention in the last few years, along with novel theoretical insights into convergence properties of upper-confidence-bound (UCB) style algorithms Liu et al. (2021); Kong et al. (2022); Zhang et al. (2022).
@@
The two-sided matching problem involves agents on two sides of a market who have preferences for each other but cannot communicate explicitly. The goal is to create a matching process that ensures stability, where no pairs of agents would rather be matched with each other over their current match. The existence of such matchings was famously demonstrated constructively in the Gale-Shapley algorithm Gale and Shapley (1962) which is structured around one side of the market proposing and the other side choosing whether to accept proposals. This theory has been applied to various markets, such as matching medical students to residencies Roth and Peranson (1999) and students to schools Abdulkadiroğlu et al. (2005), with the assumption that agents know their own preferences. Considerable interest has also been shown in the AI community regarding two-sided matching in the presence of various constraints, such as diversity constraints Aziz et al. (2022); Benabbou et al. (2018). A version of the problem, called two-sided bandits, was introduced in Das and Kamenica (2005), where agents engage in repeated matching without prior knowledge of preferences, and the impact of the matching mechanism on convergence to stability was studied. Recent work has also focused on computing the probability that a matching is stable or finding the matching with the highest probability of being stable under different models of preference uncertainty Aziz et al. (2020).
@@
The exploration-exploitation dilemma that characterizes bandit problems is further complicated in two-sided bandits by uncertainty on not just the values of the “arms” of the bandit, but also uncertainty on whether the arms will accept or reject your attempt to pull them (your proposal). This additional uncertainty arises from the multi-agent nature of the problem; instead of there being just one player, there are multiple players competing for the arms, and the arms also have preferences associated with the players. A popular solution choice for this type of problem involves the explore-then-commit style of algorithms Pagare and Ghosh (2023); Zhang and Fang (2024). This flavor of solution includes a first phase of “learning” preferences, and a second phase using what has been learned in the prior one to commit to a choice. This involves spending some dedicated time trying to figure out the preferences of the agents. In this paper we are more interested in approaches where learning happens simultaneously with the matching process. One such algorithm, CA-UCB Liu et al. (2021), assumes that at each round all the agents on the receiving side (henceforth arms) get to view all their proposals before deciding which one to accept. The algorithm provably converges to stability under the assumptions that (1) all arms have complete knowledge of their preferences, and (2) these preferences are common knowledge, so the proposers (henceforth players) also know them. The algorithm is a variation of the well-known family of UCB algorithms. Complementary to this, Hosseini et al. (2024) take a stability-centric view on the problem: using the arm-proposing version of the Gale–Shapley matching procedure, they show higher likelihood of stable outcomes and provide Probably Approximately Correct (PAC) bounds on the sample complexity of reaching stability. They do this by employing an action-elimination strategy to rule out clearly inferior arms when an agent faces multiple options, concentrating on pairs most likely to appear in a stable matching.
Contributions:
@@
We study the more general case where neither the players nor the arms start off with knowledge of their own preferences. This problem is significantly more complex than the case with one-sided uncertainty and common knowledge of the other side’s preferences. We approach the problem in two steps.
First, we relax the common knowledge assumption and assume only that arms are aware of their own preferences. We design an algorithm (Optimistic CA-UCB or OCA-UCB) that provably converges to a stable matching in this setting, using an extension of the concept of “plausible sets” (arms that an player might be able to “win” in the next time step) from the CA-UCB algorithm Liu et al. (2021). Our work differs from other recent papers that also do not assume common knowledge; those rely on either explicit communication Zhang et al. (2022) or implicit communication through carefully orchestrated “collisions” Zhang and Fang (2024).
Second, we introduce the Probabilistic Conflict Avoiding - Simultaneous Choice Algorithm (PCA-SCA) for the case with uncertainty on both sides. PCA-SCA leverages an additional optimistic estimate for the probability of winning each contentious proposal. We prove convergence to stability using UCB-style reward estimates. The framework can also accommodate other methods such as Thompson Sampling. Experimentally, we demonstrate convergence to stability under both UCB and Thompson-based reward estimates and evaluate how the convergence speed scales with the number of agents and the degree of preference heterogeneity.
2 Setting
@@
There are $N$ proposers (henceforth players) and $K$ proposees (henceforth arms). To ensure that each player can be matched, we assume that $N\leq K$ Liu et al. (2020, 2021); Basu et al. (2021); Sankararaman et al. (2021). The set of players is denoted by $\left\{p_{i}\right\}_{i=1}^{N}$ and the set of arms by $\left\{a_{k}\right\}_{k=1}^{K}$. For each $p_{i}$, there is a distinct (unknown) mean reward $\mu_{i}^{k}$ associated with each arm $a_{k}$ with unit variance. If $\mu_{i}^{k}>\mu_{i}^{j}$, we say that $p_{i}$ prefers $a_{k}$ over $a_{j}$ denoted by $a_{k}\succ_{p_{i}}a_{j}$. Similarly, the arms also have distinct mean rewards associated with each player. The arms may or may not have this information available to them a priori depending on the setting in consideration.
@@
We use the classic notion of a stable matching: a bipartite matching from the group of players to the group of arms such that no player and arm prefer each other over their current matching i.e $\nexists(p_{i},m_{i}(t))$ such that $p_{i}$ prefers an arm $a_{j}\neq m_{i}(t)$ and $a_{j}$ also prefers $p_{i}$ over its current match. There may be more than one stable match. Our primary focus in this work is on convergence to stability. Another quantity we study is maximum Player Pessimal Regret. We use the Gale-Shapley algorithm with arms as proposers in order to calculate the player-pessimal stable match Gale and Shapley (1962). Denote the reward received by player $p_{i}$ in a player pessimal match by $X_{i}$. Then, we define maximum player-regret at each time-step as $R(t)=\underset{i\in N}{\max}\left[X_{i}-X_{i}^{m(t)}\right]$ which is the maximum difference (over players) between the reward a player received in its current match and the reward it would have received in a stable matching that player-pessimal.
@@
The APCK model has been studied and a decentralized solution approach is described in the Conflict Avoiding Upper Confidence Bound (CA-UCB) algorithm Liu et al. (2021). Our main contribution is to detail decentralized approaches to solutions for Scenarios 1 and 2. We begin with a quick recap of CA-UCB in the APCK model.
APCK Model and CA-UCB
@@
The APCK model (Scenario 0) was studied in a recent paper Liu et al. (2020). Subsequently, the CA-UCB algorithm was proposed as a decentralized solution Liu et al. (2021). CA-UCB finds a stable match in the APCK model by avoiding conflicts using the notion of a “plausible set.” The plausible set for a player $p_{i}$ at time-step $t$ is defined as $S_{i}(t):=\{a_{k}:p_{i}\succ_{a_{k}}p_{j},\text{ where }\bar{A}_{j}(t-1)=a_{k}\}$ i.e. the set of all arms $a_{k}$ such that at time-step (t1)(t-1) another player $p_{j}\neq p_{i}$ successfully pulled the arm, but $p_{i}$ knows that $a_{k}$ prefers $p_{i}$ over $p_{j}$. This set also includes the arms the player successfully pulled (if any) and all the arms that were unmatched at the previous time-step. This makes it so that $p_{i}$ only attempts arms where there is a possibility of it successfully pulling the arm, avoiding conflicts. Then, $p_{i}$ attempts to pull the arm with the highest UCB value among $S_{i}(t)$ . The UCB value tracked by $p_{i}$ for an arm $a_{k}$ is calculated as $\text{UCB}_{i}^{k}=\hat{\mu}_{i}^{k}+\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t)}}$, where $\hat{\mu}_{i}^{k}$ is the empirical mean reward tracked by $p_{i}$ for $a_{k}$ up to time $t$, and $n_{i}^{k}(t)$ is the total number of samples that is used to calculate that empirical mean.
3 APKP Model and OCA-UCB
CA-UCB requires both players and arms to know arms’ preferences in order to form $S_{i}(t)$. When players lack this information, they must learn their rankings in arms’ preferences and and infer plausible sets. Recent decentralized methods address this: Etesami and Srikant Etesami and Srikant (2024) maintain and update proposal-arm probability distributions via accept/reject feedback, yielding stable matchings with logarithmic regret in hierarchical markets; Shah et al. Shah et al. (2024) employ “mood” states (content, discontent, watchful) to guide proposals based on utility feedback, converging to the proposer-optimal stable match without centralized coordination. We introduce OCA-UCB (Optimistic CA-UCB), a simpler approach that exploits arms’ self-knowledge and needs at most one “collision” per player to learn arm preferences.
Following the general structure of CA-UCB, each player initializes $\mathrm{UCB}_{i}^{k}=\infty$. Additionally, for each arm $a_{k}$, player $p_{i}$ maintains two sets, $O^{k_{h}}_{(i)}=\{p_{j}\mid p_{j}\succ_{a_{k}}p_{i}\}$ and $O^{k_{l}}_{(i)}=\{p_{j}\mid p_{i}\succ_{a_{k}}p_{j}\}$ i.e. the set of players $a_{k}$ ranks above and below $p_{i}$ in $a_{k}$’s preference order. These start optimistically as $O^{k_{h}}_{(i)}=\emptyset$ and $O^{k_{l}}_{(i)}=\{p_{j}\}_{j\neq i}$, i.e. each player assumes it is every arm’s top choice. Notice that to update plausible sets, $p_{i}$ requires only knowing, when $\bar{A}_{j}(t-1)=a_{k}$, which of $p_{i}$ or $p_{j}$ is preferred by $a_{k}$. Because the initial estimates are optimistic, all arms are plausible and conflicts become inevitable. If at time $t$ players $p_{i}$ and $p_{j}$ conflict over $a_{k}$ and $p_{j}$ wins, then $O^{k_{h}}_{(i)}\leftarrow O^{k_{h}}_{(i)}\cup\{p_{j}\}$ and $O^{k_{l}}_{(i)}\leftarrow O^{k_{l}}_{(i)}\setminus\{p_{j}\}$. Thereafter $p_{i}$ excludes $p_{j}$ when forming $S_{i}(t)$. Apart from this update, the mechanism is identical to CA-UCB. Eventually, each player’s sets recover the arms’ true orderings. We prove using Theorem 3 (stated below) that, using belief updates as mentioned above preserves CA-UCB’s guarantee of convergence. The proof is detailed in Appendix B.